Assuming that a table saw blade has a diameter of 12 inches and spins at 3450 revolutions per minute(RPMs), determine how fast the tip of the teeth of the saw blade are traveling in miles per hour.
Answer by Alan3354(69443)

Assuming that a table saw blade has a diameter of 12 inches and spins at 3450 revolutions per minute(RPMs), determine how fast the tip of the teeth of the saw blade are traveling in miles per hour.
------------
v = 12*pi*3450 in/min
v = 3450*pi ft/min
v = 3450*pi*60/5280 mi/hr
v = 3450*pi/88 mi/hr
v =~ 123 mph